# Geometric Disentanglement of Text Embeddings for Subject-Consistent Text-to-Image Generation using A Single Prompt

### v2

@@ 1 Introduction
In recent years, text-to-image diffusion models have achieved significant progress, enabling users to generate diverse, high-quality images from natural language descriptions [11, 13]. Despite their advances, such models often struggle to maintain consistent depictions of subjects across a sequence of images, a capability crucial for visual storytelling.
A dominant approach to this challenge is personalization [3, 14], where a text-to-image model is fine-tuned on user-provided images of a target subject. While effective, personalization requires time-consuming per-subject optimization and is prone to overfitting, where maintaining subject identity comes at the cost of text alignment. Alternatively, other methods incorporate image conditioning by training models with external image encoders [4, 17, 18], but these approaches demand significant computational costs.
To avoid such computational overhead, several recent works have investigated training-free strategies. StoryDiffusion [19] and ConsiStory [15] modify the self-attention mechanism of a pre-trained diffusion model to enable feature sharing across images generated in a batch to ensure subjects remain consistent. However, these methods require substantial memory to store and reuse shared features during generation. Apart from these approaches, instead of feeding the model separate text prompts for each scene, 1Prompt1Story [7] concatenates all scene descriptions into a single prompt that begins with an identity description specifying the subject’s visual attributes, followed by frame descriptions for each scene, leveraging language models’ contextual understanding capability to maintain subject consistency through shared context. Although this approach has shown plausible results, it suffers from semantic leakage, where token embeddings across frames become entangled, leading to text misalignment (see Fig. 1 top).
@@ 2.1 Consistent Text-to-Image Generation from A Single Prompt
Instead of providing the model with different text inputs for each frame, the single-prompt paradigm constructs a single text input $P=[P_{0};P_{1};...;P_{N}]$ for the model that starts with an initial identity prompt $P_{0}$, which describes the target subject’s visual attributes, followed by a series of frame prompts $P_{i}$ that specify different scenarios for each frame, implicitly maintaining identity consistency through shared context. This effect stems from the self-attention mechanism within Transformer-based text encoders [16], which allows semantic information to flow from earlier to later parts of the sequence. While this property helps maintain identity consistency as all the subsequent frame prompts $P_{i}$ attend to the prefixed identity prompt $P_{0}$, it also causes the semantics of individual frame prompts to become intertwined within the concatenated prompt embedding (see Fig. 2).
@@ 2.2 Orthogonal Semantic Subspace Projection
To mitigate semantic entanglement, we leverage the structured semantic geometry of the CLIP [12] embedding space, whose locally linear properties enable the systematic manipulation of concepts through vector operations [5, 8, 9]. Specifically, we project the text embedding matrix onto the semantic subspace of an undesired concept and subtract this projection to suppress its influence.
@@ 3.2 Evaluation Metrics
To evaluate text alignment, we compute the average CLIP-Score [6] between each generated image and its corresponding prompt, denoted as CLIP-T. For subject consistency, we measure image similarity using DreamSim [2], a validated proxy for human judgment of visual similarity, along with CLIP-I, defined as the cosine similarity between image embeddings, and DINO [1, 10], a feature similarity metric based on DINO embeddings of the generated images.

### v3

@@ 1 Introduction
In recent years, text-to-image diffusion models have achieved significant progress, enabling users to generate diverse, high-quality images from natural language descriptions [podell2023sdxlimprovinglatentdiffusion, rombach2021highresolution]. Despite their advances, such models often struggle to maintain consistent depictions of subjects across a sequence of images, a capability crucial for visual storytelling.
A dominant approach to this challenge is personalization [gal2022imageworthwordpersonalizing, ruiz2023dreamboothfinetuningtexttoimage], where a text-to-image model is fine-tuned on user-provided images of a target subject. While effective, personalization requires time-consuming per-subject optimization and is prone to overfitting, where maintaining subject identity comes at the cost of text alignment. Alternatively, other methods incorporate image conditioning by training models with external image encoders [gal2023encoderbaseddomaintuningfast, wei2023eliteencodingvisualconcepts, ye2023ipadaptertextcompatibleimage], but these approaches demand significant computational costs.
To avoid such computational overhead, several recent works have investigated training-free strategies. StoryDiffusion [zhou2024storydiffusion] and ConsiStory [tewel2024training] modify the self-attention mechanism of a pre-trained diffusion model to enable feature sharing across images generated in a batch to ensure subjects remain consistent. However, these methods require substantial memory to store and reuse shared features during generation. Apart from these approaches, instead of feeding the model separate text prompts for each scene, 1Prompt1Story [liu2025onepromptonestoryfreelunchconsistenttexttoimage] concatenates all scene descriptions into a single prompt that begins with an identity description specifying the subject’s visual attributes, followed by frame descriptions for each scene, leveraging language models’ contextual understanding capability to maintain subject consistency through shared context. Although this approach has shown plausible results, it suffers from semantic leakage, where token embeddings across frames become entangled, leading to text misalignment (see Fig. 1 top).
@@ 2.1 Consistent Text-to-Image Generation from A Single Prompt
Instead of providing the model with different text inputs for each frame, the single-prompt paradigm constructs a single text input $P=[P_{0};P_{1};...;P_{N}]$ for the model that starts with an initial identity prompt $P_{0}$, which describes the target subject’s visual attributes, followed by a series of frame prompts $P_{i}$ that specify different scenarios for each frame, implicitly maintaining identity consistency through shared context. This effect stems from the self-attention mechanism within Transformer-based text encoders [vaswani2023attentionneed], which allows semantic information to flow from earlier to later parts of the sequence. While this property helps maintain identity consistency as all the subsequent frame prompts $P_{i}$ attend to the prefixed identity prompt $P_{0}$, it also causes the semantics of individual frame prompts to become intertwined within the concatenated prompt embedding (see Fig. 2).
@@ 2.2 Orthogonal Semantic Subspace Projection
To mitigate semantic entanglement, we leverage the structured semantic geometry of the CLIP [radford2021learningtransferablevisualmodels] embedding space, whose locally linear properties enable the systematic manipulation of concepts through vector operations [grand2018semanticprojectionrecoveringhuman, mikolov2013distributedrepresentationswordsphrases, mikolov-etal-2013-linguistic]. Specifically, we project the text embedding matrix onto the semantic subspace of an undesired concept and subtract this projection to suppress its influence.
@@ 3.2 Evaluation Metrics
To evaluate text alignment, we compute the average CLIP-Score [hessel2022clipscorereferencefreeevaluationmetric] between each generated image and its corresponding prompt, denoted as CLIP-T. For subject consistency, we measure image similarity using DreamSim [fu2023dreamsim], a validated proxy for human judgment of visual similarity, along with CLIP-I, defined as the cosine similarity between image embeddings, and DINO [caron2021emergingpropertiesselfsupervisedvision, oquab2023dinov2], a feature similarity metric based on DINO embeddings of the generated images.